In an elevator, a loose 0.4kg square ceiling tile is only held in place by the friction force from the four surrounding tiles (The coefficient of static friction between tiles is 0.500). (a) If the normal force from each tile is 2.00 N what acceleration of the elevator will cause the tile to fall on and unsuspecting physics professor's head? (b) In what direction must the elevator be accelerating for this to happen?
Answer #1

a] Maximum static friction from each tile = 0.5*(2+2+2+2) = 4 N

Let the acceleration for falling be a, we take upward as positive direction.

Friction - mg = ma

4 - 0.4*9.8 = 0.4*a

a = 10 - 9.8 = 0.2 m/s^2 answer

b] direction is upward as we got positive value in last part .